OpenStudy (anonymous):

Consider the function f(x) = x^3+4x^2+kx-4 the remainder from f(x)/(x+2) is twice the remainder from f(x) /(x-2) determine the the value of k

OpenStudy (raden):

remainder of f(x)/(x+2), is f(-2) and remainder of f(x)/(x-2), is f(2). f(x) = x^3+4x^2+kx-4, thus f(-2)= (-2)^3 + 4(-2)^2 + k(-2) - 4 f(2)= (2)^3 + 4(2)^2 + k(2) - 4 so, if f(-2)=2f(2) then (-2)^3 + 4(-2)^2 + k(-2) - 4 = 2( (2)^3 + 4(2)^2 + k(2) - 4 ) simplify,, and solve for k ! goodluck...